OpenStudy (anonymous):

factor: 9x2 -(2y-1)2

OpenStudy (anonymous):

\[9x^2 -(2y-1)^2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[9x^2-(2y-1)^2 =(3x+2y-1)(3x-2y+1)\] Using \[a^2-b^2 = (a+b)(a-b)\]